Tohban Report 2012-04-04

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 28 Mar 2012
End Date: 4 Apr 2012
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Martin Fivian
List all reports


Solar Activity

This was a very quiet week, with no flares above C level. The past few days have been especially quiet, with no flares above B level.

How many GOES flares occurred?

Flares above B, C, M, X class were     16    7     0     0

And how many of these are listed in the RHESSI flare list?

Flares above B, C, M, X class were     12     5     0     0

And how many had EXCELLENT coverage?

Flares above B, C, M, X class were      8     4     0     0

Memory Management

The SSR accidentally reached 100% due to a special commanding event (see Operations section below). The SSR filled completely at 1115 UT on April 4, 2012. At 1545 UT, the SSR write pointer was moved back to 1000 UT, so RHESSI will not have data for the interval 1000 - 1545 on this day. SSR level is currently at 70% and dropping on each pass. Weilheim passes are also being added to deal with this issue, and we should be back to normal within a week or so. Since solar activity is so low, it was decided not to take any further (drastic) steps to reduce the SSR fill level.

RHESSI Operations

April 3, 22:26:42 UT: Commands were sent to increase the SAS data rate for a 7-hour period coordinated with an SDO roll maneuver for limb observations. Although this commanding was expected to only cause a small increase in the SSR fill, it instead caused the SSR to fill completely. This may have been due to a miscalculation in the amount of data collected; this is being looked into and no misbehavior of the ADP is suspected. We guess that the first 3 orbits during the SDO roll maneuver were acquired; this data will be examined for quality once it is received on the ground.

March 30, 17:53:27 UT: The front fast event threshold for detector 2 was raised by 0x10, from 0x60 to 0x70. This was done in order to save livetime that was being eaten up by excess fast rates in G2 due to light leak.

Data Gaps

There were no gaps in monitor rates other than the 5-6 hour gap noted above due to the SSR filling.

Detector Issues

The (front fast) threshold increase in detector 2 appears to have quieted that detector. Light leakage is still causing increased fast events in this detector, but the livetime is not so drastically affected. David pointed out that there is a tradeoff to consider, in that a higher threshold (probably) means that we cannot use G2 for spectroscopy at low energies, but a lower threshold will mean that the detector saturates in large flares. Which of these is the lesser evil is open to debate! It was also noted that the detector response needs to be updated both to account for this change and for the changes at/since the last anneal. In order to figure out the effect of the raised threshold level on the detector's efficiency, spectra need to be compared for G2 and some other detectors.

Detector 8 is showing elevated front slow rates coming out of the SAA. It then slowly quiets down over the course of the orbit. The effect on the livetime is small, but it is not yet clear whether this occurs at energies that can affect spectroscopy. This will be investigated.

Personal tools